# CAML: Collaborative Auxiliary Modality Learning for Multi-Agent Systems

### v2

@@ Abstract
Multi-modal learning has become a crucial technique for improving the performance of machine learning applications across domains such as autonomous driving, robotics, and perception systems. However, in certain scenarios, particularly in resource-constrained environments, some modalities available during training may be absent during inference. While existing frameworks effectively utilize multiple data sources during training and enable inference with reduced modalities, they are primarily designed for single-agent settings. This poses a critical limitation in dynamic environments such as connected autonomous vehicles (CAV), where incomplete data coverage can lead to decision-making blind spots. Conversely, some works explore multi-agent collaboration but without addressing missing modality at test time. To overcome these limitations, we propose Collaborative Auxiliary Modality Learning (CAML), a novel multi-modal multi-agent framework that enables agents to collaborate and share multi-modal data during training, while allowing inference with reduced modalities during testing. Experimental results in collaborative decision-making for CAV in accident-prone scenarios demonstrate that CAML achieves up to a ${\bf 58.1}\%$ improvement in accident detection. Additionally, we validate CAML on real-world aerial-ground robot data for collaborative semantic segmentation, achieving up to a ${\bf 10.6}\%$ improvement in mIoU.
@@ 1 Introduction
Multi-modal learning has become an essential approach in a wide range of machine learning applications, particularly in areas such as autonomous driving [6, 37, 7], robotics [27, 20], and perception systems [42, 2], where the availability of multiple data sources (e.g., RGB images, LiDAR, radar) improves model performance by providing complementary information. However, these multi-modal systems often incur increased computational overhead and latency at inference time. More critically, in real-world conditions, some modalities may be unavailable, or too costly to acquire at test time, which challenges the reliability and efficiency of such systems.
Recent work in machine learning [16, 36, 9, 10, 28] has sought to address these problems by enabling models to leverage additional modalities during training while supporting inference with fewer modalities. Such approaches reduce computational costs and improve robustness in conditions where some sensors may be unavailable at inference time. Shen et al. [32] formalized these tasks under the framework of Auxiliary Modality Learning (AML), which effectively reduces the dependency on expensive or unreliable modalities.
@@
To address these limitations, we propose Collaborative Auxiliary Modality Learning (CAML), a novel framework for multi-modal, multi-agent systems. CAML enables agents to collaborate and share multi-modal data during training, while supporting inference with reduced modalities at test time, as illustrated in Fig. 1. This collaborative learning process enhances scene understanding and data coverage by allowing agents to compensate for each other’s blind spots, leading to more informed predictions and decision-making. To ensure effective deployment in scenarios where only partial modalities are available (e.g., due to sensor failures or resource constraints), CAML utilizes knowledge distillation [15] to transfer knowledge from a teacher model, trained with full-modality inputs, to a student model that learns to perform effectively with reduced modalities at inference time. This distillation process embeds full-modality knowledge into the student model, preserving the benefits of collaborative training and ensuring robust inference even when some modalities are absent. For instance, in autonomous driving, multiple vehicles can share sensor information such as LiDAR and RGB images during training to build more robust representations, while during runtime testing, each vehicle performs inference using only RGB images.
Unlike prior work that either focuses on multi-agent collaboration but without addressing missing modality at test time, or explores AML ideas solely in single-agent settings, CAML unifies these concepts. Overall, this work offers the following key contributions:
@@
  • We validate CAML through experiments in collaborative decision-making for connected autonomous driving in accident-prone scenarios, and collaborative semantic segmentation for real-world data of aerial-ground robots. CAML achieves up to ${\bf 58.1}\%$ improvement in accident detection for autonomous driving, and up to ${\bf 10.6}\%$ improvement for more accurate semantic segmentation.
@@ 2 Related Work
Collaboration in multi-agent systems has been widely studied across fields such as autonomous driving and robotics. In autonomous driving, prior research has explored various strategies, including spatio-temporal graph neural networks [8], LiDAR-based end-to-end systems [5], decentralized cooperative lane-changing [26] and game-theoretic models [13]. In robotics, Mandi et al. [25] presented a hierarchical multi-robot collaboration approach using large language models, while Zhou et al. [40] proposed a perception framework for multi-robot systems built on graph neural networks. A review of multi-robot systems in search and rescue operations was provided by [30], and Bae et al. [1] developed a reinforcement learning (RL) method for multi-robot path planning. Additionally, various communication mechanisms, such as Who2com [22], When2com [21], and Where2comm [17], have been created to optimize agent interactions.
Despite these advancements, existing multi-agent collaboration frameworks remain limited by their focus on specific tasks and the assumption that agents will have consistent access to the same data modalities during both training and testing, an assumption that may not hold in real-world applications. To address these gaps, our framework CAML enables agents to collaborate during training by sharing multi-modal data, but at test time, each agent performs inference using reduced modality. This reduces the dependency on certain modalities for deployment, while still allowing agents to leverage additional information during training to enhance overall performance and robustness.
@@
Auxiliary Modality Learning (AML) [32] has emerged as an effective solution to reduce computational costs and the amount of input data required for inference. By utilizing auxiliary modalities during training, AML minimizes reliance on those modalities at inference time. For example, Hoffman et al. [16] introduced a method that incorporates depth images during training to enhance test-time RGB-only detection models. Similarly, Wang et al. [36] proposed PM-GANs to learn a full-modal representation using data from partial modalities, while Garcia et al. [9, 10] developed approaches that use depth and RGB videos during training but rely solely on RGB data for testing. Piasco et al. [28] created a localization system that predicts depth maps from RGB query images at test time. Building on these works, Shen et al. [32] formalized the AML framework, systematically classifying auxiliary modality types and AML architectures.
However, existing AML frameworks are typically designed for single-agent settings, failing to exploit the potential benefits of multi-agent collaboration for improving multi-modal learning. In contrast, CAML allows agents to collaboratively learn richer multi-modal representations. By sharing complementary information, this approach enhances overall data coverage and mitigates information loss during reduced-modality inference, which is particularly important when an individual agent’s observations are incomplete due to occlusions or limited sensor range.
@@
Knowledge distillation (KD) [15] is a widely used technique in many domains to reduce computation by transferring knowledge from a large, complex model (teacher) to a simpler model (student). In computer vision, Gou et al. [11] provided a comprehensive survey of KD applications, while Beyer et al. [3] conducted an empirical investigation to develop a robust and effective recipe for making large-scale models more practical. Additionally, Tung and Mori [35] introduced a KD loss function that aligns the training of a student network with input pairs producing similar activation in the teacher network. In natural language processing, Xu et al. [38] reviewed the applications of KD in LLMs, while Sun et al. [33] proposed a Patient KD method to compress larger models into lightweight counterparts that maintain effectiveness. Hahn and Choi [12] also suggested a KD approach that leverages the soft target probabilities of the training model to train other neural networks. In autonomous driving, Lan and Tian [19] presented an approach for visual detection, Cho et al. [4], Sautier et al. [31] used KD for 3D object detection.
Notice that existing KD mostly distills knowledge from a larger model to a smaller one to reduce computation, Shen et al. [32] aimed to design a cross-modal learning approach using KD to utilize the hidden information from auxiliary modalities within the AML framework. But AML is limited by the scope of a single-agent paradigm. In contrast, we leverage KD within multi-agent settings, where the teacher models are trained with access to shared multi-modal data (e.g., RGB and LiDAR) from multiple agents. By distilling this collaborative knowledge into each agent’s reduced modality (e.g., RGB), CAML enables robust inference during deployment, even with fewer modalities. This collaborative distillation process enhances each agent’s performance by providing richer, complementary knowledge from the collaborative training phase.
@@ 3 Collaborative Auxiliary Modality Learning
In single-agent frameworks such as AML [32], the missing modalities during testing are referred to as auxiliary modalities, while those that remain available are called the main modality. In contrast, in our framework CAML, each agent can process a different number of modalities during training and different agents can have different main modalities and auxiliary modalities. There is no correlation between the number of agents and the number of modalities.
Figure 2: Overview of the CAML Approach Pipeline. The teacher model (top) aggregates and shares multi-modal embeddings across agents to make predictions using the full set of modalities. In contrast, the student model (bottom) processes a subset of modalities per agent and shares them to form a multi-modal embedding. Through knowledge distillation from the teacher, the student learns to produce robust predictions despite missing modalities, enabling effective inference during deployment. In the teacher model, the set of agents is denoted as ${\mathcal{A}}_{\text{train}}=\{{\mathcal{A}}_{1},{\mathcal{A}}_{2},\ldots,{\mathcal{A}}_{N}\}$. The set of modalities is denoted as ${\mathcal{I}}_{\text{train}}$. The observations of all agents are denoted as $X=\{x_{1},x_{2},\ldots,x_{N}\}$, where $x_{i}^{k}$ is the observation acquired by the $i$-th agent ${\mathcal{A}}_{i}\in{\mathcal{A}}_{\text{train}}$ for the $k$-th modality. In the student model, the set of agents is denoted as ${\mathcal{A}}_{\text{test}}=\{{\mathcal{A}}_{1},{\mathcal{A}}_{2},\ldots,{\mathcal{A}}_{M}\}$. The set of modalities is denoted as ${\mathcal{I}}_{\text{test}}$, which is a subset of ${\mathcal{I}}_{\text{train}}$. The set of agents that have access to the $j$-th modality ${\mathcal{I}}_{j}\in{\mathcal{I}}_{\text{test}}$ is denoted as ${\mathcal{A}}_{\text{test}}^{{\mathcal{I}}_{j}}$, and the number of agents in this set is given by $|{\mathcal{A}}_{\text{test}}^{{\mathcal{I}}_{j}}|=M_{j}$, with $x_{M_{j}}^{j}$ represents the observation acquired by the $M_{j}$-th agent ${\mathcal{A}}_{M_{j}}\in{\mathcal{A}}_{\text{test}}$ for the $j$-th modality.
@@
After designing the CAML framework, we present a detailed complexity analysis. If the system is centralized, all collaborative agents share their data with one designated ego agent for centralized processing. Each of the $N-1$ collaborative agents performs its local computation independently, with a time complexity of $O(T_{c})$ and a space complexity of $O(S_{c})$, where $T_{c}$ represents the time required for local computation, and $S_{c}$ is the associated space. Thus, the total computation time and space complexities for all collaborative agents are $O(T_{c}(N-1))$ and $O(S_{c}(N-1))$, respectively. For simplicity, assuming each communication from one collaborative agent to the ego agent consumes $O(D)$ time complexity and $O(M)$ space complexity, where $D$ is the time required for communication and $M$ is the corresponding space. Therefore, the total communication time and space complexities for gathering information at the ego agent are $O(D(N-1))$ and $O(M(N-1))$, respectively. Then the ego agent aggregates the received data, running a model, having a time and space complexity $O(T_{e})$ and $O(S_{e})$, where $T_{e}$ and $S_{e}$ represent the time and space required for the ego agent’s computation. So the total time and space complexities are $O(T_{c}(N-1)+D(N-1)+T_{e})$ and $O(S_{c}(N-1)+M(N-1)+S_{e})$, respectively.
If the system is decentralized, each agent performs its local computation and shares information with nearby agents. For simplicity, let the local computation for a single agent has a time complexity of $O(T)$, where $T$ is the time required for local computation. Assume that communication from one agent to another agent requires $O(D)$ time complexity and $O(M)$ space complexity, where $D$ represents the time of communication between two agents, and $M$ denotes the space required for such communication. For $N$ agents, the total computation time complexity is $O(NT)$. In the worst case, each agent share data with all other agents, this can result in $O(N^{2}D)$ for pairwise sharing. So the total time complexity is $O(NT+N^{2}D)$. For space complexity, the storage requirement for all agents is $O(NS)$, where $S$ is the space needed per agent. Communication between agents adds an additional complexity of $O(N^{2}M)$. So the total space complexity is $O(NS+N^{2}M)$. In the typical case, if each agent communicates with only other $k$ agents ($k\ll N$) rather than all $N-1$ agents. The total time and space complexities become $O(NT+NkD)$ and $O(NS+NkM)$, respectively.
@@ 4.1 Collaborative Decision-Making
The dataset is generated using the AUTOCASTSIM benchmark [29], which features three complex and accident-prone traffic scenarios designed for CAV. These scenarios are characterized by limited sensor coverage and obstructed views, as illustrated in Fig. 6. Following prior works COOPERNAUT [5] and STGN [8], we employ an expert agent with full access to the traffic scene information to collect data. This expert agent is deployed in the environment to generate the data trails, where each trail captures RGB images and LiDAR point clouds from both the ego vehicle and collaborative vehicles, along with the control actions of the ego vehicle. In line with the setup used in STGN [8], the ego vehicle can have up to three collaborative vehicles at each timestep, provided they are within a 150-meter radius. Each trail represents a unique instantiation of the traffic scenario, with differences arising from the randomized scenario configurations. For further dataset details, please refer to Appendix A.2.1.
@@
After data collection, we train the models using Behavior Cloning (BC), following the setup established by prior works COOPERNAUT [5] and STGN [8]. For each vehicle, both RGB and LiDAR data are used during training, while only RGB data is used during testing in CAML. Please refer to Appendix A.2.3 for more details on the BC procedure. All experiments are conducted across four repeated runs, and we report the mean and standard deviation of the results.
For processing RGB data, we first resize the image to $224\times 224$ and use ResNet-18 [14] as the encoder to extract a feature map for each vehicle. We then apply self-attention on the feature map to dynamically compute the importance of features at different locations. After the self-attention, we apply three convolution layers with each followed by a ReLU activation. Finally, we obtain a 256-d feature representation after passing through a fully connected layer. To aggregate RGB feature embeddings from connected vehicles to the ego vehicle, we use the cross-attention mechanism. For processing the LiDAR data, we use the Point Transformer as the encoder for each vehicle and utilize the COOPERNAUT [5] model to aggregate LiDAR feature embeddings between connected vehicles. Then we concatenate the final RGB and LiDAR embeddings for the ego vehicle’s decision-making, with a three-layer MLP as the prediction module to output the action.
@@
We implement the following baselines for comparison: (1) AML [32]: In the AML setting, the ego vehicle operates independently without collaboration with other vehicles (non-collaborative). Both RGB and LiDAR data are available during training for the vehicle, while only RGB data is available during testing. (2) COOPERNAUT [5]: A collaborative method that processes LiDAR data during both training and testing. It employs the Point Transformer [39] as the backbone, encoding raw 3D point clouds into keypoints. (3) STGN [8]: A collaborative approach that utilizes spatial-temporal graph networks for decision-making, with RGBD data used for both training and testing.
@@
Additionally, how does CAML compare with other approaches that have access to more modalities during testing? We evaluate CAML against STGN [8] under multi-agent settings. CAML uses only RGB data during testing but STGN uses both RGB and depth data. Despite with fewer modality, CAML achieves comparable, and in some cases superior performance while relying solely on RGB data, as illustrated in Fig. 4. Notably, CAML exceeds the mean ADR of STGN by ${\bf 10.9\%}$ in the left-turn scenario, demonstrating that our model can enhance driving safety even when constrained to fewer modalities. This further underscores the strength of CAML, which effectively leverages LiDAR data as an auxiliary modality during training to boost performance. The fact that CAML matches or exceeds the performance of a model that uses more data at test time highlights the modality efficacy of our approach.
@@
We compare the performance of our approach with COOPERNAUT [5] and STGN [8] under single-agent settings, with results shown in Fig. 4. CAML outperforms the other baselines across all scenarios, for both ADR and EIR metrics. This demonstrates that even with a single agent during testing, CAML remains highly effective, by utilizing the multi-agent collaboration and auxiliary modalities provided by the teacher model during training.
Table 1: Communication and inference efficiency comparison. CAML can operate more efficiently with lower communication overhead and faster inference than both [5] and [8].
@@
COOPERNAUT [5] 65.5 90.0
STGN [8] 4.9 18.5
@@
We evaluate the communication and inference efficiency of CAML compared to other cooperative approaches, COOPERNAUT [5] and STGN [8]. To assess communication bandwidth, we follow STGN [8] and use the shared package size (PS) metric, which measures the size of data exchanged between connected vehicles, serving as an indicator of communication efficiency in collaborative decision-making. For inference efficiency, we measure latency per batch, which reflects the average time required to process a batch during inference. The evaluation results are summarized in Table 1. As shown, CAML achieves lower PS and latency compared to other approaches, highlighting its superior communication and inference efficiency.
@@
In the ablation studies, we compare the performance of CAML with two other baselines: the fully-equipped teacher model, and a multi-agent BC model trained and tested using only RGB images, without knowledge distillation (KD). As shown in Table 2, while CAML shows a slight performance drop compared to the teacher, due to missing modalities at inference, it still achieves strong and robust performance. This highlights the effectiveness of our method in handling modality reduction scenarios, making it especially suitable for resource-constrained environments. Furthermore, CAML consistently outperforms the RGB-only BC baseline. The results demonstrate the advantage of incorporating multi-modal learning into multi-agent systems. Specifically, leveraging additional modalities during training through knowledge distillation allows CAML to encode richer information into the student model, resulting in improved performance at test time.

### v3

@@ Abstract
Multi-modal learning has emerged as a key technique for improving performance across domains such as autonomous driving, robotics, and reasoning. However, in certain scenarios, particularly in resource-constrained environments, some modalities available during training may be absent during inference. While existing frameworks effectively utilize multiple data sources during training and enable inference with reduced modalities, they are primarily designed for single-agent settings. This poses a critical limitation in dynamic environments such as connected autonomous vehicles (CAV), where incomplete data coverage can lead to decision-making blind spots. Conversely, some works explore multi-agent collaboration but without addressing missing modality at test time. To overcome these limitations, we propose Collaborative Auxiliary Modality Learning (CAML), a novel multi-modal multi-agent framework that enables agents to collaborate and share multi-modal data during training, while allowing inference with reduced modalities during testing. Experimental results in collaborative decision-making for CAV in accident-prone scenarios demonstrate that CAML achieves up to a $\bf 58.1$$\%$ improvement in accident detection. Additionally, we validate CAML on real-world aerial-ground robot data for collaborative semantic segmentation, achieving up to a $\bf 10.6$$\%$ improvement in mIoU.
@@ 1 Introduction
Multi-modal learning has become an essential approach in a wide range of machine learning systems, particularly in areas such as autonomous driving [el2019rgb, xiao2020multimodal, gao2018object, liu2025mmcd], robotics [noda2014multimodal, lee2020making, liu2025imrl], and reasoning [liu2025vogue, li2025self, liu2025stable], where the availability of multiple data sources (e.g., RGB, LiDAR, radar, text) improves model performance by providing complementary information. However, these multi-modal systems often incur increased computational overhead and latency at inference time. More critically, in real-world conditions, some modalities may be unavailable, or too costly to acquire at test time, which challenges the reliability and efficiency of such systems.
Recent work in machine learning [hoffman2016learning, wang2018pm, garcia2018modality, garcia2019learning, piasco2021improving] has sought to address these problems by enabling models to leverage additional modalities during training while supporting inference with fewer modalities. Such approaches reduce computational costs and improve robustness in conditions where some sensors may be unavailable at inference time. shen2023auxiliary formalized these tasks under the framework of Auxiliary Modality Learning (AML), which effectively reduces the dependency on expensive or unreliable modalities.
@@
To address these limitations, we propose Collaborative Auxiliary Modality Learning (CAML), a novel framework for multi-modal, multi-agent systems. CAML enables agents to collaborate and share multi-modal data during training while supporting inference under reduced-modality conditions at test time, as illustrated in Fig. 1. This collaborative learning paradigm enhances scene understanding and data coverage by allowing agents to compensate for each other’s blind spots, leading to more informed and coordinated decision-making. To ensure robust deployment in scenarios with missing modalities (e.g., due to sensor failures or resource constraints), CAML employs knowledge distillation (KD) [hinton2015distilling] to transfer knowledge from a teacher model trained with full-modality inputs to a student model that performs effectively with limited modalities. By enabling cross-modal and cross-agent knowledge transfer, CAML extends traditional KD to handle partial observability and support team-level coordination. The distillation process embeds rich, full-modality knowledge into the student model, preserving the benefits of collaborative training and ensuring reliable inference even when some modalities are unavailable. For example, in autonomous driving, multiple vehicles can share LiDAR and RGB data during training to build robust shared representations, while at deployment each vehicle performs inference using only RGB inputs.
Unlike prior work that either focuses on multi-agent collaboration but without addressing missing modality at test time, or explores AML ideas solely in single-agent settings, CAML unifies these concepts. To the best of our knowledge, this work is the first to propose a flexible and principled framework for multi-modal multi-agent systems where limited modalities are available for inference. Overall, this work offers the following key contributions:
@@
  • We validate CAML through extensive experiments in collaborative decision-making for connected autonomous driving in accident-prone scenarios, and collaborative semantic segmentation for real-world data of aerial-ground robots. CAML achieves up to $\bf 58.1$$\%$ improvement in accident detection for autonomous driving, and up to $\bf 10.6$$\%$ improvement for more accurate semantic segmentation.
@@ 2 Related Work
Collaboration in multi-agent systems has been widely studied across fields such as autonomous driving and robotics. In autonomous driving, prior research has explored various strategies, including spatio-temporal graph neural networks [gao2024collaborative], LiDAR-based end-to-end systems [cui2022coopernaut], decentralized cooperative lane-changing [nie2016decentralized] and game-theoretic models [hang2021decision]. In robotics, mandi2024roco presented a hierarchical multi-robot collaboration approach using large language models, while zhou2022multi proposed a perception framework for multi-robot systems built on graph neural networks. A review of multi-robot systems in search and rescue operations was provided by [queralta2020collaborative], and bae2019multi developed a reinforcement learning (RL) method for multi-robot path planning. Additionally, various communication mechanisms, such as Who2com [liu2020who2com], When2com [liu2020when2com], and Where2comm [hu2022where2comm], have been created to optimize agent interactions.
Despite these advancements, existing multi-agent collaboration frameworks remain limited by their focus on specific tasks and the assumption that agents will have consistent access to the same data modalities during both training and testing, an assumption that may not hold in real-world applications. To address these gaps, CAML enables agents to collaborate during training by sharing multi-modal data, but at test time, each agent performs inference using reduced modality. This reduces the dependency on certain modalities for deployment, while still allowing agents to leverage additional information during training to enhance overall performance and robustness.
@@
Auxiliary Modality Learning (AML) [shen2023auxiliary] has emerged as an effective solution to reduce computational costs and the amount of input data required for inference. By utilizing auxiliary modalities during training, AML minimizes reliance on those modalities at inference time. For example, hoffman2016learning introduced a method that incorporates depth images during training to enhance test-time RGB-only detection models. Similarly, wang2018pm proposed PM-GANs to learn a full-modal representation using data from partial modalities, while garcia2018modality, garcia2019learning developed approaches that use depth and RGB videos during training but rely solely on RGB data for testing. piasco2021improving created a localization system that predicts depth maps from RGB query images at test time. Building on these works, shen2023auxiliary formalized the AML framework, systematically classifying auxiliary modality types and AML architectures.
However, existing AML frameworks are typically designed for single-agent settings, failing to exploit the potential benefits of multi-agent collaboration for improving multi-modal learning. Transitioning from single-agent to multi-agent learning is fundamentally non-trivial, as multi-agent systems introduce unique challenges such as dynamic team compositions and team-level decision-making. CAML effectively addresses these challenges and allows agents to collaboratively learn richer multi-modal representations. By sharing complementary information, CAML enhances overall data coverage and mitigates information loss during reduced-modality inference, which is particularly important when an individual agent’s observations are incomplete due to occlusions or limited sensor range.
@@
Knowledge distillation (KD) [hinton2015distilling] is a widely used technique in many domains to reduce computation by transferring knowledge from a large, complex model (teacher) to a simpler model (student). In computer vision, gou2021knowledge provided a comprehensive survey of KD applications, while beyer2022knowledge conducted an empirical investigation to develop a robust and effective recipe for making large-scale models more practical. Additionally, tung2019similarity introduced a KD loss function that aligns the training of a student network with input pairs producing similar activation in the teacher network. In natural language processing, xu2024survey reviewed the applications of KD in LLMs, while sun2019patient proposed a Patient KD method to compress larger models into lightweight counterparts that maintain effectiveness. hahn2019self suggested a self-distillation approach that leverages the soft target probabilities of a model to train other neural networks, while liu2025aukt explored KD under uncertainty. In autonomous driving, lan2022instance presented an approach for visual detection, cho2023itkd, sautier2022image used KD for 3D object detection.
Notice that existing KD mostly distills knowledge from a larger model to a smaller one to reduce computation, shen2023auxiliary aimed to design a cross-modal learning approach using KD to utilize the hidden information from auxiliary modalities within the AML framework. But AML is limited by the scope of a single-agent paradigm. In contrast, we leverage KD within multi-agent settings, where the teacher models are trained with access to shared multi-modal data (e.g., RGB and LiDAR) from multiple agents. By distilling this collaborative knowledge into each agent’s reduced modality (e.g., RGB), CAML enables robust inference during deployment, even with fewer modalities. This collaborative distillation process enhances each agent’s performance by providing richer, complementary knowledge from the collaborative training phase.
@@ 3 Collaborative Auxiliary Modality Learning
In single-agent frameworks such as AML [shen2023auxiliary], the missing modalities during testing are referred to as auxiliary modalities, while those that remain available are called the main modality. In contrast, in our framework CAML, each agent can process a different number of modalities during training and different agents can have different main modalities and auxiliary modalities. There is no correlation between the number of agents and the number of modalities.
Figure 2: Overview of the Pipeline of CAML. The teacher model (top) aggregates and shares multi-modal embeddings across agents to make predictions using the full set of modalities. In contrast, the student model (bottom) processes a subset of modalities per agent and shares them to form a multi-modal embedding. Through knowledge distillation from the teacher, the student learns to produce robust predictions despite missing modalities, enabling effective inference during deployment. In the teacher model, the set of agents is denoted as ${\mathcal{A}}_{\text{train}}=\{{\mathcal{A}}_{1},{\mathcal{A}}_{2},\ldots,{\mathcal{A}}_{N}\}$. The set of modalities is denoted as ${\mathcal{I}}_{\text{train}}$. The observations of all agents are denoted as $X=\{x_{1},x_{2},\ldots,x_{N}\}$, where $x_{i}^{k}$ is the observation acquired by the $i$-th agent ${\mathcal{A}}_{i}\in{\mathcal{A}}_{\text{train}}$ for the $k$-th modality. In the student model, the set of agents is denoted as ${\mathcal{A}}_{\text{test}}=\{{\mathcal{A}}_{1},{\mathcal{A}}_{2},\ldots,{\mathcal{A}}_{M}\}$. The set of modalities is denoted as ${\mathcal{I}}_{\text{test}}$, which is a subset of ${\mathcal{I}}_{\text{train}}$. The set of agents that have access to the $j$-th modality ${\mathcal{I}}_{j}\in{\mathcal{I}}_{\text{test}}$ is denoted as ${\mathcal{A}}_{\text{test}}^{{\mathcal{I}}_{j}}$, and the number of agents in this set is given by $|{\mathcal{A}}_{\text{test}}^{{\mathcal{I}}_{j}}|=M_{j}$, with $x_{M_{j}}^{j}$ represents the observation acquired by the $M_{j}$-th agent ${\mathcal{A}}_{M_{j}}\in{\mathcal{A}}_{\text{test}}$ for the $j$-th modality.
Problem Formulation.
@@
After designing the CAML framework, we present a detailed complexity analysis. If the system is centralized, all collaborative agents share their data with one designated ego agent for centralized processing. Each of the N1N-1 collaborative agents performs its local computation independently, with a time complexity of $O(T_{c})$ and a space complexity of $O(S_{c})$, where $T_{c}$ represents the time required for local computation, and $S_{c}$ is the associated space. Thus, the total computation time and space complexities for all collaborative agents are $O(T_{c}(N-1))$ and $O(S_{c}(N-1))$, respectively. For simplicity, assuming each communication from one collaborative agent to the ego agent consumes $O(D)$ time complexity and $O(M)$ space complexity, where $D$ is the time required for communication and $M$ is the corresponding space. Therefore, the total communication time and space complexities for gathering information at the ego agent are O(D(N1))O(D(N-1)) and O(M(N1))O(M(N-1)), respectively. Then the ego agent aggregates the received data, running a model, having a time and space complexity $O(T_{e})$ and $O(S_{e})$, where $T_{e}$ and $S_{e}$ represent the time and space required for the ego agent’s computation. So the total time and space complexities are $O(T_{c}(N-1)+D(N-1)+T_{e})$ and $O(S_{c}(N-1)+M(N-1)+S_{e})$, respectively.
If the system is decentralized, each agent performs its local computation and shares information with nearby agents. For simplicity, let the local computation for a single agent has a time complexity of $O(T)$, where $T$ is the time required for local computation. Assume that communication from one agent to another agent requires $O(D)$ time complexity and $O(M)$ space complexity, where $D$ represents the time of communication between two agents, and $M$ denotes the space required for such communication. For $N$ agents, the total computation time complexity is $O(NT)$. In the worst case, each agent share data with all other agents, this can result in $O(N^{2}D)$ for pairwise sharing. So the total time complexity is $O(NT+N^{2}D)$. For space complexity, the storage requirement for all agents is $O(NS)$, where $S$ is the space needed per agent. Communication between agents adds an additional complexity of $O(N^{2}M)$. So the total space complexity is $O(NS+N^{2}M)$. In the typical case, if each agent communicates with only other $k$ agents ($k\ll N$) rather than all N1N-1 agents. The total time and space complexities become $O(NT+NkD)$ and $O(NS+NkM)$, respectively.
@@ 4.1 Collaborative Decision-Making
The dataset is generated using the AUTOCASTSIM benchmark [qiu2021autocast], which features three complex and accident-prone traffic scenarios designed for CAV. These scenarios are characterized by limited sensor coverage and obstructed views, as illustrated in Fig. 6. Following prior works COOPERNAUT [cui2022coopernaut] and STGN [gao2024collaborative], we employ an expert agent with full access to the traffic scene information to collect data. This expert agent is deployed in the environment to generate the data trails, where each trail captures RGB images and LiDAR point clouds from both the ego vehicle and collaborative vehicles, along with the control actions of the ego vehicle. In line with the setup used in STGN [gao2024collaborative], the ego vehicle can have up to three collaborative vehicles at each timestep, provided they are within a 150-meter radius. Each trail represents a unique instantiation of the traffic scenario, with differences arising from the randomized scenario configurations. For further dataset details, please refer to Appendix A.2.1.
@@
After data collection, we train the models using Behavior Cloning (BC) [liu2024adaptive, bhaskar2024lava], following the setup established by prior works COOPERNAUT [cui2022coopernaut] and STGN [gao2024collaborative]. For each vehicle, both RGB and LiDAR data are used during training, while only RGB data is used during testing in CAML. Please refer to Appendix A.2.3 for more details on the BC procedure. All experiments are conducted across four repeated runs, and we report the mean and standard deviation of the results.
For processing RGB data, we first resize the image to $224\times 224$ and use ResNet-18 [he2016deep] as the encoder to extract a feature map for each vehicle. We then apply self-attention on the feature map to dynamically compute the importance of features at different locations. After the self-attention, we apply three convolution layers with each followed by a ReLU activation. Finally, we obtain a 256-d feature representation after passing through a fully connected layer. To aggregate RGB feature embeddings from connected vehicles to the ego vehicle, we use the cross-attention mechanism. For processing the LiDAR data, we use the Point Transformer as the encoder for each vehicle and utilize the COOPERNAUT [cui2022coopernaut] model to aggregate LiDAR feature embeddings between connected vehicles. Then we concatenate the final RGB and LiDAR embeddings for the ego vehicle’s decision-making, with a three-layer MLP as the prediction module to output the action.
@@
We implement the following baselines for comparison: (1) AML [shen2023auxiliary]: In the AML setting, the ego vehicle operates independently without collaboration with other vehicles (non-collaborative). Both RGB and LiDAR data are available during training for the vehicle, while only RGB data is available during testing. (2) COOPERNAUT [cui2022coopernaut]: A collaborative method that processes LiDAR data during both training and testing. It employs the Point Transformer [zhao2021point] as the backbone, encoding raw 3D point clouds into keypoints. (3) STGN [gao2024collaborative]: A collaborative approach that utilizes spatial-temporal graph networks for decision-making, with RGBD data used for both training and testing.
@@
Additionally, how does CAML compare with other approaches that have access to more modalities during testing? We evaluate CAML against STGN [gao2024collaborative] under multi-agent settings. CAML uses only RGB data during testing but STGN uses both RGB and depth data. Despite with fewer modality, CAML achieves comparable, and in some cases superior performance while relying solely on RGB data, as illustrated in Fig. 4. Notably, CAML exceeds the mean ADR of STGN by $\bf 10.9\%$} in the left-turn scenario, demonstrating that our model can enhance driving safety even when constrained to fewer modalities. This further underscores the strength of CAML, which effectively leverages LiDAR data as an auxiliary modality during training to boost performance. The fact that CAML matches or exceeds the performance of a model that uses more data at test time highlights the modality efficacy of our approach.
@@
We compare the performance of our approach with COOPERNAUT [cui2022coopernaut] and STGN [gao2024collaborative] under single-agent settings, with results shown in Fig. 4. CAML outperforms the other baselines across all scenarios, for both ADR and EIR metrics. This demonstrates that even with a single agent during testing, CAML remains highly effective, by utilizing the multi-agent collaboration and auxiliary modalities provided by the teacher model during training.
Table 1: Communication and inference efficiency comparison. CAML can operate more efficiently with lower communication overhead and faster inference than both [cui2022coopernaut] and [gao2024collaborative].
@@
COOPERNAUT [cui2022coopernaut] 65.5 90.0
STGN [gao2024collaborative] 4.9 18.5
@@
We evaluate the communication and inference efficiency of CAML compared to other cooperative approaches, COOPERNAUT [cui2022coopernaut] and STGN [gao2024collaborative]. To assess communication bandwidth, we follow STGN [gao2024collaborative] and use the shared package size (PS) metric, which measures the size of data exchanged between connected vehicles, serving as an indicator of communication efficiency in collaborative decision-making. For inference efficiency, we measure latency per batch, which reflects the average time required to process a batch during inference. The evaluation results are summarized in Table 1. As shown, CAML achieves lower PS and latency compared to other approaches, highlighting its superior communication and inference efficiency.
@@
In the ablation studies, we first compare the performance of CAML with two other baselines: the fully-equipped teacher model, and a multi-agent BC model trained and tested using only RGB images, without knowledge distillation (KD). As shown in Table 2, while CAML shows a slight performance drop compared to the teacher, due to missing modalities at inference, it still achieves strong and robust performance. This highlights the effectiveness of our method in handling modality reduction scenarios, making it especially suitable for resource-constrained environments. Furthermore, CAML consistently outperforms the RGB-only BC baseline. The results demonstrate the advantage of incorporating multi-modal learning into multi-agent systems. Specifically, leveraging additional modalities during training through knowledge distillation allows CAML to encode richer information into the student model, resulting in improved performance at test time.